OpenStudy (anonymous):

1. The following is a geometric sequence. -18, 12, -8, 5, …. True or false?

OpenStudy (anonymous):

Do you know how to find first and common ratio from the geometric sequence given??

OpenStudy (anonymous):

first term I mean..

OpenStudy (klimenkov):

Try to check the condition \[\frac {12} {-18} = \frac {-8} {12}=\frac {5} {-8}\] If it is true, you have geometric sequence.

OpenStudy (anonymous):

The first two equal each other but the last one isn''t, so this is false?

OpenStudy (klimenkov):

You are right.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

geometric sequence general formula is \[a_{n}=a_{1}r^{n-1}\] find the common ratio 12/-18=-2/3 -8/12=-2/3 5/8=5/8 since the common ratio is not same so it is not geometric